OpenStudy (anonymous):

Please help: Solve f(x)= x^2-4x/ radical x

OpenStudy (anonymous):

\[f(x)=\frac{ x^2-4x }{\sqrt{x} }=\frac{ x \left( x-4 \right) }{\sqrt{x} }=\sqrt{x}\left( x-4 \right)\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw